slope

What is the equation of the line that passes through the point (-4,2) and has a slope of -\frac{5}{4}?
Answer #1
Question:
Write an equation of the line thathas the given slope and passes theough the given point. Usepoint-slope and slope intercept form:
equation:
m = 5/2, P IS THEPOINT (X1,Y1)=(-4,2)
X1=-4
Y1=2
FORMULA IS
Y-Y1=M(X-X1)
Y-2=(5/2)[X-(-4)] = (5X/2)+5*4/2 = (5X/2) + 10
Y= 5X/2 +10+2
Y=(5X/2)+12 OR
Y=2.5X+12 OR
Y=(5/2)X+12
Answer #2
m = 5/2, (-4,2)

y - y1 = m(x - x1)    now plug it in
y - 2 = 5/2(x + 4)
y - 2 = 5x/2 + 10    plus 2 to bothside
y = 5x/2 +12     <---- equation
now plug in the point and test the equation
(-4, 2)   
2 = 5(-4)/2 + 12
2 = -10 + 12
2 = 2
IT WORKED